OpenStudy (anonymous):

What is the area, in square units, of an isosceles right triangle with a hypotenuse of 20 units?

OpenStudy (jhannybean):



OpenStudy (anonymous):

20^2 = 2x^2 area of triangle should be (x^2)/2

OpenStudy (jhannybean):

I love this drawing tool.

OpenStudy (anonymous):

how do you change color and size?

OpenStudy (jhannybean):

Its a chrome extension.

OpenStudy (anonymous):

how to get from 20 to x lol

OpenStudy (anonymous):

20^2 = 20*20 = ?

OpenStudy (anonymous):

oh pythagorean theorem?

OpenStudy (anonymous):

I think I worked it out

OpenStudy (jhannybean):

First figure out what x is equal to. \[20^2 = x^2+x^2 \]\[20^2 = 2x^2\]\[\color{red}{x^2}=\frac{20^2}{2}\]\[A_{\triangle} = \frac{1}{2}b\cdot h =\frac{1}{2} x \cdot x =\frac{1}{2}\color{red}{x^2} \]

OpenStudy (anonymous):

thx

OpenStudy (jhannybean):

No problem.